OpenStudy (anonymous):

a(b + c),if a = 3, b = -2, and c = 5

OpenStudy (cj49):

substitute the values 3(-2+5)=9

OpenStudy (anonymous):

Substitute You have 3(-2 + 5) Step 1 : 3 + -2 = 1 Step 2 : 3 + 5 = 8 Step 3 : 1 + 8 = 9 Your answer = 9

OpenStudy (phi):

you can do this 2 ways. the first way is pedmas. Do the parens first: (-2+5) which is the same as 5-2, simplifies to 3 you now have 3*3 which is 9 the other way is to "distribute" the 3 in 3(-2+5) that means multiply 3 times each number inside the parens 3*-2 + 3*5 simplify to get -6 + 15 and 15-6 = 9 just like before.